OpenStudy (anonymous):

HELPPPP!!!!! using the method of your choice to solve the system below write solution as an ordered pair. x+2=5 y=2

OpenStudy (anonymous):

x+2 = 5 x = 5-2 x = 3 The ordered pair would be (3,2) because x is 3 and y is 2